Why the income and interest sensitivities of the demand for real balances affect the slope of LM curve?


1
Expert's answer
2021-11-04T10:23:07-0400

IS curve showing inverse relation between interest and output. IS curve showing goods market in equillibrium. While the LM curve showing direct relation interest and output. LM curve showing money and asset market in equilibrium.

The slope of the LM curve affected by both income and interest sensitivity for showing this

"LM \space equation\\L=Ky-Hi\\\frac{M}{P}=Ky-Hi"

Here K is showing income sensitivity and H is showing interest sensitivity. For knowing slope of the LM curve change this equation as i is dependent variable:

"\frac{M}{P}=Ky-Hi\\i=(\frac{K}{H})y-(\frac{1}{H})\frac{M}{P}"

It means the slope of LM equation is income sensitivity divided by interest sensitivity.K is direct proportionate with the slope and H is indirect. I.e., if K increases then the slope increases and LM curve is steeper. If H increases then the slope decreases and the LM curve is flatter.

The slope of the LM curve increases with increase in income sensitivity and the slope of the LM curve decreases with increase in interest sensitivity